# Non-Markovian Reduced Models to Unravel Transitions in Non-equilibrium Systems

### v1

@@ Abstract
This work proposes a general framework for capturing noise-driven transitions in spatially extended non-equilibrium systems and explains the emergence of coherent patterns beyond the instability onset. The framework relies on stochastic parameterization formulas to reduce the complexity of the original equations while capturing the key effects of unresolved scales. This approach works for both Gaussian and Lévy-type noise.
Our stochastic parameterization formulas offer two key advantages. First, they can approximate stochastic invariant manifolds when these manifolds exist. Second, even when such manifolds break down, our formulas can be adapted through a simple optimization of its constitutive parameters. This allows us to handle scenarios with weak time-scale separation where the system has undergone multiple transitions, resulting in large-amplitude solutions not captured by invariant manifold or other time-scale separation methods.
@@ I Introduction
The emergence of coherent, complex patterns is ubiquitous in many spatially extended non-equilibrium systems; see e.g. [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. Mechanisms to explain the emergence of these patterns include the development of instability saturated by nonlinear effects whose calculations can be conducted typically near the onset of linear instability; see e.g. [11, 12]. Historically, hydrodynamic systems have been commonly used as prototypes to study instabilities out of equilibrium in spatially extended systems, from both an experimental and a theoretical point of view [13, 3].
@@
However, for non-equilibrium systems away from the onset of linear instability, the emergence of coherent patterns does not fit within this instability/nonlinear saturation theory as the reduction principle of the fast modes onto the slow ones breaks down [11] and calls for new reduction techniques to explain emergence. Furthermore, in the presence of random fluctuations, reduced equations under the form of deterministic ODEs are inherently incapable to capture phenomena like noise-induced transitions. Noise can have unexpected outcomes such as altering the sequence of transitions, possibly suppressing instabilities [14], the onset of turbulence [15] or, to the opposite, exciting transitions [16, 7]. Examples in which noise-induced transitions have an important role include the generation of convective rolls [17, 18], electroconvection in nematic liquid crystals [16, 19], certain oceanic flows [20, 21], and climate dynamics [22, 23, 24, 25].
@@
To address these limitations, the framework of optimal parameterizing manifolds (OPMs) introduced in [26, 27, 28] for forced-dissipative systems offers a powerful solution. This framework allows for the efficient derivation of parameterizations away from the onset of instability, achieved through continuous deformations of near-onset parameterizations and optimized by data-driven minimization of a discrepancy metric measuring the defect of parameterization of the unresolved variables by the resolved ones. At the core of the approach are backward-forward systems derived from the original model [27, 28], which allow for efficient parameterizations that do not suffer the restrictions of invariant manifolds (spectral gap condition [27, Eq. (2.18)]) after data-driven optimization.
@@
In this article, we carry over this variational approach to the case of spatially extended non-equilibrium systems within the framework of stochastic partial differential equations (SPDEs). In particular, we extend the parameterization formulas obtained in [29, 30, 31] for SPDEs driven by multiplicative noise (parameter noise [18, 7]), to SPDEs driven, beyond this onset, by more realistic, spatio-temporal noise either of Gaussian or of Lévy type. In particular, our framework allows for dealing with the important case of cutoff scales larger than the scales forced stochastically.
For this case, the stochastic parameterizations we derive give rise to reduced systems taking the form of stochastic differential equations (SDEs) with non-Markovian, path-dependent coefficients depending on the noise’s past. Non-Markovian coefficients, like those derived here, have been encountered in the reduction of stochastic systems near a change of stability, where we can rely on the natural separation of time-scales. There, several methods have indeed pointed out the occurrence of these coefficients in the derivation of simpler equations to describe the evolution of the dominant pattern. These include approaches based on stochastic normal forms [32, 33, 34], stochastic invariant manifold approximations [35, 29, 30, 31], or multiscale methods in which the order of the noise variance is comparable to the order of the distance from the change of stability [36, 37, 38, 39, 31]. Although the forcing noise is Gaussian, these non-Markovian terms can obey non-Gaussian statistics in the case of multiplicative noise; see [31, Appendix A].
These exogenous memory effects, stemming from the forcing noise’s history, are distinct from endogenous memory effects encountered in the reduction of nonlinear autonomous systems [40, 41, 42]. The latter, predicted by the Mori-Zwanzig (MZ) theory, are functionals of the past of the resolved state variables. Endogenous memory effects often appear when the validity of deterministic invariant manifolds or the conditional expectation in the MZ expansion, breaks down [43, 27, 44, 42]. Due to their independence from resolved variables, exogenous memory effects offer computational advantages, allowing efficient computations without integral quadratures (see Section III.2).
Building upon [27, 28], we demonstrate that optimizing the memory length of these non-Markovian coefficients efficiently captures the interplay between nonlinearities and noise in the original SPDE. This enables us to derive reduced SDEs that faithfully replicate the SPDE’s stochastic dynamics for regimes beyond the instability onset, where the system has undergone multiple transitions away from its basic state and multiple coherent patterns coexist.
Our parameterization framework is tested against the stochastic Allen-Cahn equation (sACE), a powerful tool for modeling non-equilibrium phase transitions in various scientific fields [1, 3, 45, 46]. However, analyzing such spatially extended systems to understand the interplay between noise and nonlinearities driving stochastic transitions remains a challenge, especially when multiple coherent patterns coexist. Understanding these stochastic transitions holds significant importance for a broader perspective, impacting multiple scientific disciplines and offering valuable insights into real-world phenomena; see e.g. [47] and references therein.
Despite its importance in numerous physical applications, analyzing the reduction problem to capture transitions in the sACE has received limited attention, with existing works focusing primarily on scenarios near instability onset and low noise intensity (e.g., [48, 49, 50, 51, 52]). Solving this problem for regimes outside instability onset could significantly impact the application of emerging rare event algorithms to more complex, high-dimensional system (e.g., [53, 54, 55, 21, 47]). In that respect, our approach addresses the limitations of standard reduction methods, by enabling the derivation of novel stochastic normal forms that emphasize the crucial role of non-Markovian coefficients in specific circumstances.
@@
This work sheds light on a key example of such circumstances, applicable not only to the sACE but also to the reduction of many other stochastic models. It occurs when the cutoff scale, defining the retained large-scale dynamics, is chosen to be larger than the forcing scale, which characterizes the characteristic length of the imposed noise. In this case, any traditional deterministic reduced models at the cutoff scale based on e.g. Galerkin methods or deterministic parameterizations, filter out the underlying noise’s effects. This leaves them blind to the subtle fluctuations driving the system’s behavior and leading eventually to stochastic transitions. In contrast, our stochastic parameterization approach sees right through this filter. It tracks how the noise, acting in the “unseen” part of the system, interacts with the resolved modes through the aforementioned, intricate non-Markovian coefficients.
Our parameterization formulas extend beyond the specific case of sACE. They are applicable to a wider range of SPDEs with cubic nonlinearities, particularly those exhibiting degenerate forcing, where the cutoff scale is chosen just above the forcing scale. While existing multiscale methods, which assume strong time-scale separation and proximity to instability onset, can provide for such degenerate forcing scenarios, rigorous error estimates [56], these methods often fall short in physically relevant scenarios with weak time-scale separation. Our approach offers a way to overcome these limitations from a practical standpoint.
We show in Section V below, how the nonlinear and stochastic terms in our parameterizations, when trained over a single path, enable us to derive reduced equations that retain a remarkable predictive power even in presence of weak-time scale separation. In that respect, the derived reduced equations that are only stochastically driven by the non-Markovian coefficients in our formalism, are not only able to accurately emulate ensemble statistics (averages across many realizations) but also anticipate the system’s typical metastable states as well as those reached through the rarest transition paths.
Our stochastic parameterization approach is also robust against various noise types, including Lévy noise. To showcase this versatility, we apply it to another important category of spatially extended non-equilibrium systems: jump-driven SPDEs whose unperturbed dynamics exhibit an S-shaped continuum of steady states. These solution curves are indicative of multistability, tipping phenomena and critical transitions [57, 58], as well as hysteresis loops [59, 60], and are found in diverse applications. These include combustion theory [61, 62], plasma physics [63, 64], ecosystems [9], neurosciences [65], Earth’s energy balance models [66, 67], and ocean dynamics [68, 69]. More recently, critical transitions and tipping phenomena have played a crucial role in understanding the climate crisis, with evidence from observations and general circulation models [70, 71, 72], conceptual network approaches [73], and theoretical studies [42].
SPDEs driven by jump processes, have recently garnered significant attention due to their specific property in the way noise-induced transitions between coexisting attractors take place. Unlike the Gaussian case, these transitions occur through rare, individually impactful jumps [74]. Jump processes have demonstrated their versatility in a diverse range of fields. Notably, they offer valuable insights into multistable behaviors observed in complex systems [75, 76, 77, 78]. Furthermore, jump processes have proven effective in modeling real-world phenomena such as Dansgaard-Oeschger events in paleoclimate records [79], chaotic transport in rotating flows [80], missing physics in precipitation models [81, 82], and tropical convection dynamics [83, 84].
While SPDEs driven by jump processes have seen increasing applications in various fields, efficient techniques for reducing their complexity are still under development (e.g. [85, 86, 87]). This highlights the potential of our proposed approach. In that respect, we demonstrate the flexibility of the stochastic parameterization framework of Section III by showcasing in Section VI its effectiveness in capturing the intricate interplay between jump noise and nonlinear dynamics within the reduced equations.
@@ II.1 Spatially extended non-equilibrium systems
$A=-L+B,$ (2)
where $L$ is sectorial with domain $D(L)\subset H$ which is compactly and densely embedded in $H$. We assume also that $-L$ is stable, while $B$ is a low-order perturbation of $L$, i.e. $B$ is a bounded linear operator such that $B:D(L^{\alpha})\rightarrow H$ for some $\alpha$ in $[0,1)$. We refer to [90, Sec. 1.4] for an intuitive presentation of fractional power of an operator and characterization of the so-called operator domain $D(L^{\alpha})$. In practice, the choice of $\alpha$ should match the loss of regularity effects caused by the nonlinear terms so that
$G\colon D(L^{\alpha})\rightarrow H,$
@@
with $p>k\geq 2$, and $G_{k}$ denoting the leading-order operator (on $D(L^{\alpha})$) in the Taylor expansion of $G$, near the origin. A broad class of spatially extended stochastic equations from physics can be recasted into this framework (see [92, 88, 29, 93, 31] for examples), as well as time-delay systems subject to stochastic disturbances [82, 94].
@@
where the $\bm{e}_{j}$ are eigenmodes of $A$, the $\eta_{t}^{j}$ denote either a finite family of mutually independent jump processes, or mutually independent Brownian motions, $W_{t}^{j}$, over its probability space $(\Omega,\mathbb{P},\mathcal{F})$ with $\mathbb{P}$ denoting the canonical Wiener measure [34, Appendix A.3].
@@
$H_{\mathfrak{c}}\oplus H_{\mathfrak{s}}=H.$ (6)
@@
$\Pi_{\mathfrak{c}}:H\rightarrow H_{\mathfrak{c}},\textrm{ and }\Pi_{\mathfrak{s}}:H\rightarrow H_{\mathfrak{s}}.$ (7)
@@
Our goal is to provide reduced models able to capture with fidelity the large-scale dynamics without having to resolve the small scales.
@@
Within our working assumptions, the spectrum $\sigma(A)$ consists only of isolated eigenvalues with finite multiplicities. This combined with the sectorial property of $A$ implies that there are at most finitely many eigenvalues with a given real part. The sectorial property of $A$ also implies that the real part of the spectrum, $\Re\sigma(A)$, is bounded above (see also [95, Thm. II.4.18]). These two properties of $\Re\sigma(A)$ allow us in turn to label elements in $\sigma(A)$ according to the lexicographical order which we adopt throughout this article:
@@
$\Re\lambda_{n}>\Re\lambda_{n^{\prime}},$ (9)
@@
$\Re\lambda_{n}=\Re\lambda_{n^{\prime}},\quad\text{ and }\quad\Im\lambda_{n}\geq\Im\lambda_{n^{\prime}}.$ (10)
This way, we can rely on a simple labeling of the eigenvalues/eigenmodes by positive integers to organize the resolved and unresolved scales and the corresponding parameterization formulas derived hereafter. In practice, when the spatial domain is 2D or 3D, it is usually physically more intuitive to label the eigenelements by wave vectors. The parameterization formulas presented below can be easily recast within this convention.

### v2

@@ Abstract
This work proposes a general framework for analyzing noise-driven transitions in spatially extended non-equilibrium systems and explains the emergence of coherent patterns beyond the instability onset. The framework relies on stochastic parameterization formulas to reduce the complexity of the original equations while preserving the essential dynamical effects of unresolved scales. The approach is flexible and operates for both Gaussian noise and non-Gaussian noise with jumps.
Our stochastic parameterization formulas offer two key advantages. First, they can approximate stochastic invariant manifolds when these manifolds exist. Second, even when such manifolds break down, our formulas can be adapted through a simple optimization of its constitutive parameters. This allows us to handle scenarios with weak time-scale separation where the system has undergone multiple transitions, resulting in large-amplitude solutions not captured by invariant manifolds or other time-scale separation methods.
@@ I Introduction
Non-equilibrium systems are irreversible systems characterized by a continuous flow of energy that is driven by external forces or internal fluctuations. There are many different types of non-equilibrium systems, and they can be found in a wide variety of fields, including physics, chemistry, biology, and engineering. Non-equilibrium systems can exhibit complex behavior, including self-organization, pattern formation, and chaos. These complex behaviors arise from the interplay of nonlinear dynamics and statistical fluctuations. The emergence of coherent, complex patterns is ubiquitous in many spatially extended non-equilibrium systems; see e.g. [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. Mechanisms to explain the emergence of these patterns include the development of instability saturated by nonlinear effects whose calculations can be conducted typically near the onset of linear instability; see e.g. [11, 12]. Historically, hydrodynamic systems have been commonly used as prototypes to study instabilities out of equilibrium in spatially extended systems, from both an experimental and a theoretical point of view [13, 3].
@@
However, for non-equilibrium systems away from the onset of linear instability, such as when the Reynolds number for a fluid flow increases far beyond a laminar regime, the emergence of coherent patterns does not fit within this instability/nonlinear saturation theory as the reduction principle of the fast modes onto the slow ones breaks down [11] and calls for new reduction techniques to explain emergence. Furthermore, in the presence of random fluctuations, reduced equations under the form of deterministic ODEs are inherently incapable to capture phenomena like noise-induced transitions. Noise can have unexpected outcomes such as altering the sequence of transitions, possibly suppressing instabilities [14], the onset of turbulence [15] or, to the opposite, exciting transitions [16, 7]. Examples in which noise-induced transitions have an important role include the generation of convective rolls [17, 18], electroconvection in nematic liquid crystals [16, 19], certain oceanic flows [20, 21], and climate phenomena [22, 23, 24, 25].
@@
To address these limitations, the framework of optimal parameterizing manifolds (OPMs) introduced in [26, 27, 28] for forced-dissipative systems offers a powerful solution. This framework allows for the efficient derivation of parameterizations away from the onset of instability, achieved through continuous deformations of near-onset parameterizations that are optimized by data-driven minimization. Typically, the loss function involves a natural discrepancy metric measuring the defect of parameterization of the unresolved variables by the resolved ones. The parameterizations that are optimized are derived from the governing equations by means of backward-forward systems providing useful approximations of the unstable growth saturated by the nonlinear terms, even when the spectral gap between the stable and unstable modes is small [27, 29, 28]. These approximations, once optimized, do not suffer indeed the restrictions of invariant manifolds (spectral gap condition [27, Eq. (2.18)]) allowing to handle situations with e.g. weak time-scale separation between the resolved and unresolved modes.
@@
In this article, we carry over this variational approach to the case of spatially extended non-equilibrium systems within the framework of stochastic partial differential equations (SPDEs). In particular, we extend the parameterization formulas obtained in [30, 29, 31] for SPDEs driven by multiplicative noise (parameter noise [18, 7]), to SPDEs driven, beyond this onset, by more realistic, spatio-temporal noise either of Gaussian or of non-Gaussian nature with jumps. For this class of SPDEs, our framework allows for dealing with the important case of cutoff scales larger than the scales forced stochastically.
For this forcing scenario, the stochastic parameterizations derived in this work, give rise to reduced systems taking the form of stochastic differential equations (SDEs) with non-Markovian, path-dependent coefficients depending on the noise’s past; see e.g. Section V.2 below. We mention that many time series records from real-world datasets, are known to exhibit long-term memory. This is the case of long-range memory features assumed to represent the internal variability of the climate on time scales from years to centuries [32, 33, 34, 35]. There, the surface temperature is considered as a superposition of internal variability (the response to stochastic forcing) and a forced signal which is the linear response to external forcing. The background variability may be modeled using a stochastic process with memory, or a different process that incorporates non-Gaussianity if this is considered more appropriate [36]. As will become apparent at the end of this paper, our reduction approach is adaptable to PDEs under the influence of stochastic forcing with long-range dependence; see Section IX.
In parallel and since Hasselmann’s seminal work [37], several climate models have been incorporating stochastic forcing to represent unresolved processes [38, 39, 40, 41, 42, 43]. Our approach shows that reduced models derived from these complex systems are expected to exhibit finite-range memory effects depending on the past history of the stochastic forcing, even if the latter is white in time. These exogenous memory effects, stemming from the forcing noise’s history, are distinct from endogenous memory effects encountered in the reduction of nonlinear, unforced systems [44, 45, 43]. The latter, predicted by the Mori-Zwanzig (MZ) theory, are functionals of the past of the resolved state variables. Endogenous memory effects often appear when the validity of the conditional expectation in the MZ expansion, breaks down [46, 26, 47, 43]. The exogenous memory effects dealt with in this work, have a different origin, arising as soon as the cutoff scale is larger than the scales forced stochastically.
It is worth mentioning that reduced models with non-Markovian coefficients responsible for such exogenous memory effects, have been encountered in the reduction of stochastic systems driven by white noise, albeit near a change of stability. These are obtained from reduction approaches benefiting from timescale separation such as stochastic normal forms [48, 49, 50], stochastic invariant manifold approximations [51, 30, 29, 31], or multiscale methods [52, 53, 54, 55]. Our reduction approach is not limited to timescale separation.
In that respect, our reduction framework is tested against a stochastic Allen-Cahn equation (sACE), a powerful tool for modeling non-equilibrium phase transitions in various scientific fields [1, 3, 56, 57]. The latter model is set in a parameter regime with weak timescale separation and far from the instability onset, in which the system exhibits multiple coexisting metastable states connected to the basic state through rare or typical stochastic transition paths.
We show that our resulting OPM reduced systems, when trained over a single path, are not only able to retain a remarkable predictive power to emulate ensemble statistics (correlations, power spectra, etc.) obtained as average over a large ensemble of noise realizations, but also anticipate what are the system’s typical and rare metastable states and their statistics of occurrence. The OPM reduced system’s ability to reproduce accurately such transitions is rooted in its very structure. Its coefficients are nonlinear functionals of the aforementioned non-Markovian terms (see Eq. (85) below) allowing for an accurate representation of the genuine nonlinear interactions between noise and nonlinear terms in the original sACE, which drive fluctuations in the large-mode amplitudes.
@@
We emphasize, that because the cutoff scale, defining the retained large-scale dynamics, is chosen here to be larger than the forcing scale, any deterministic reduction (e.g. Galerkin, invariant manifold, etc.) would filter out the underlying noise’s effects. This leaves them blind to the subtle fluctuations driving the system’s behavior and leading eventually to stochastic transitions. In contrast, our stochastic parameterization approach sees right through this filter. It tracks how the noise, acting in the “unseen” part of the system, interacts with the resolved modes through our reduction method’s non-Markovian coefficients.
To demonstrate the broad applicability of our reduction approach, we apply it in Section VI to another significant class of spatially extended non-equilibrium systems: jump-driven SPDEs. The unforced, nonlinear dynamics is here chosen to exhibit an S-shaped solution curve. Many systems sharing this attribute are indicative of multistability, tipping points, and hysteresis. Such behaviors are observed in various fields, including combustion theory [58, 59], plasma physics [60, 61], ecology [9], neuroscience [62], climate science [63, 64], and oceanography [65, 66]. By understanding the complex interactions between noise and nonlinearity into these phenomena, we can gain insights into critical transitions and tipping points, which are increasingly relevant to addressing global challenges like climate change [67, 68, 69, 70, 43].
Jump processes offer a powerful tool for modeling complex systems with non-smooth dynamics. They offer valuable insights into multistable behaviors [71, 72] and have been applied to various real-world phenomena, such as paleoclimate events [73, 36], chaotic transport [74], atmospheric dynamics [75, 76, 77, 78, 79, 80], and cloud physics [81, 82].
SPDEs driven by jump processes are gaining increasing attention in applications [83]. For instance, jump processes can replace (non-smooth) ”if-then” conditions in such models, enabling more efficient simulations [80]. Nevertheless, efficient reduction techniques to disentangle the jump interactions with other smoother nonlinear components of the model, are still under development [84, 85, 86]. Our stochastic parameterization framework provides a promising solution in this direction, as demonstrated in Section VI. By capturing the intricate interplay between jump noise and nonlinear dynamics within the reduced equations, our approach can significantly simplify the analysis of such complex systems.
@@ II.1 Spatially extended non-equilibrium systems
$A=-\mathcal{L}+\mathcal{B},$ (2)
where $\mathcal{L}$ is sectorial with domain $D(\mathcal{L})\subset H$ which is compactly and densely embedded in $H$. We assume also that $-\mathcal{L}$ is stable, while $\mathcal{B}$ is a low-order perturbation of $\mathcal{L}$, i.e. $\mathcal{B}$ is a bounded linear operator such that $\mathcal{B}:D(\mathcal{L}^{\alpha})\rightarrow H$ for some $\alpha$ in $[0,1)$. We refer to [89, Sec. 1.4] for an intuitive presentation of fractional power of an operator and characterization of the so-called operator domain $D(\mathcal{L}^{\alpha})$. In practice, the choice of $\alpha$ should match the loss of regularity effects caused by the nonlinear terms so that
$G\colon D(\mathcal{L}^{\alpha})\rightarrow H,$
@@
with $p>k\geq 2$, and $G_{k}$ denoting the leading-order operator (on $D(\mathcal{L}^{\alpha})$) in the Taylor expansion of $G$, near the origin. A broad class of spatially extended stochastic equations from physics can be recasted into this framework (see [91, 87, 30, 92, 31] for examples), as well as time-delay systems subject to stochastic disturbances [82, 93].
@@
where the $\bm{e}_{j}$ are eigenmodes of $A$, the $\eta_{t}^{j}$ denote either a finite family of mutually independent jump processes, or mutually independent Brownian motions, $W_{t}^{j}$, over their relevant probability space $(\Omega,\mathbb{P},\mathcal{F})$ endowed with its canonical probability measure $\mathbb{P}$ and filtration $\mathcal{F}$; see e.g. [50, Appendix A.3].
@@
$H_{\mathfrak{c}}\oplus H_{\mathfrak{f}}=H.$ (6)
@@
$\Pi_{\mathfrak{c}}:H\rightarrow H_{\mathfrak{c}},\textrm{ and }\Pi_{\mathfrak{f}}:H\rightarrow H_{\mathfrak{f}}.$ (7)
@@
Our goal is to provide a general approach to derive reduced models that preserve the essential features of the large-scale dynamics without resolving the small scales.
We present hereafter and in Section III below, the formulas of the underlying small-scale parameterizations in the Gaussian noise case. The formalism is flexible and easily adaptable to the case of non-Gaussian noises with jumps such as discussed in Sections VI and VII, below.
@@
Within our working assumptions, the spectrum $\sigma(A)$ consists only of isolated eigenvalues with finite multiplicities. This combined with the sectorial property of $A$ implies that there are at most finitely many eigenvalues with a given real part. The sectorial property of $A$ also implies that the real part of the spectrum, $\mathrm{Re}(\sigma(A))$, is bounded above (see also [94, Thm. II.4.18]). These two properties of $\mathrm{Re}(\sigma(A))$ allow us in turn to label elements in $\sigma(A)$ according to the lexicographical order which we adopt throughout this article:
@@
$\mathrm{Re}(\lambda_{n})>\mathrm{Re}(\lambda_{n^{\prime}}),$ (9)
@@
$\mathrm{Re}(\lambda_{n})=\mathrm{Re}(\lambda_{n^{\prime}}),\quad\text{ and }\quad\mathrm{Im}(\lambda_{n})\geq\mathrm{Im}(\lambda_{n^{\prime}}).$ (10)
This way, we can rely on a simple labeling of the eigenvalues/eigenmodes by positive integers to organize the resolved and unresolved scales and the corresponding parameterization formulas derived hereafter. In practice, when the spatial domain is 2D or 3D, it is usually physically more intuitive to label the eigenelements by wave vectors. The parameterization formulas presented below can be easily recast within this convention; see e.g. [95].